# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

### v1

@@ Abstract
Statistical inference with finite-sample validity for the value function of a given policy in Markov decision processes (MDPs) is crucial for ensuring the reliability of reinforcement learning. Temporal Difference (TD) learning, arguably the most widely used algorithm for policy evaluation, serves as a natural framework for this purpose. In this paper, we study the consistency properties of TD learning with Polyak-Ruppert averaging and linear function approximation, and obtain three significant improvements over existing results. First, we derive a novel sharp high-dimensional probability convergence guarantee that depends explicitly on the asymptotic variance and holds under weak conditions. We further establish refined high-dimensional Berry-Esseen bounds over the class of convex sets that guarantee faster rates than those in the literature. Finally, we propose a plug-in estimator for the asymptotic covariance matrix, designed for efficient online computation. These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function, with guaranteed finite-sample coverage. We demonstrate the applicability of our theoretical findings through numerical experiments.
@@ 1 Introduction
Statistical inference tasks, such as constructing confidence regions or simultaneous confidence intervals, are often addressed by deriving distributional theory such as central limit theorems (CLTs) for the estimator in use. Due to the high dimensionality of modern science and engineering applications, there has been a surge of interests in deriving convergence results that are valid in a finite-sample manner. In Reinforcement Learning (RL), a discipline that underpins many recent machine learning breakthroughs (Sutton and Barto, (2018); Agarwal et al., (2019)) one central question is to evaluate with confidence the quality of a given policy, measured by its value function. As RL is often modeled as decision making in Markov decision processes (MDPs), the task of statistical inference needs to accommodate the online nature of the sampling mechanism.
Temporal Difference (TD) learning (Sutton, (1988)) is arguably the most widely used algorithm designed for this purpose. TD learning, which is an instance of stochastic approximation (SA) method (Robbins and Monro, (1951)), approximates the value function of a given policy in an iterative manner. Towards understanding the non-asymptotic convergence rate of TD to the target value function, extensive recent efforts have been made (see, e.g. Nemirovski et al., (2009); Rakhlin et al., (2012); Li et al., 2024a ; Bhandari et al., (2021); Dalal et al., (2018); Khamaru et al., (2021); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018); Mou et al., (2020); Durmus et al., (2022); Patil et al., (2023); Li et al., 2024b ; Samsonov et al., (2023) and references therein) and it is known that for the tabular case, where both state and action spaces are assumed to be finite, the TD algorithm is minimax optimal (Li et al., 2024a ). Another thread of work considers the asymptotic performance of TD where it has been shown that with Polyak-Ruppert averaging (Polyak and Juditsky, (1992); Ruppert, (1998)), TD converges to a Gaussian distribution in the large sample limit (Fort, (2015); Mou et al., (2020); Li et al., (2023)). However, these two types of results remain insufficient for conducting statistical inference that is valid in the finite-sample fashion.
@@
To close this gap, in this paper, based on the TD learning algorithm, we design confidence regions and simultaneous confidence intervals for the linear coefficients of the value function when the it is approximated by a linear combination of features. To achieve this goal, we first provide a novel sharp high-dimensional probability convergence guarantee for the TD iterates, and follow up with refined high-dimensional Berry-Esseen bounds over the class of convex sets. Finally, we propose a plug-in estimator for the asymptotic covariance matrix, improving upon other alternatives proposed by recent literature (Chen et al., (2020), White and White, (2010), Hanna et al., (2017), Hao et al., (2021), Ramprasad et al., (2022)) in both computational and statistical efficiency. We summarize our main contributions in details in the following section. While we are writing this paper, we become aware of a great recent paper by Samsonov et al., (2024), which considers similar problems to ours and derives a Berry-Esseen bound for linear SA with Polyak-Ruppert averaging. It remains unclear whether their general framework gives tight results when specialized to TD learning. We offer more detailed comparisons between our results and theirs in the following sections.
1.1 Contributions
Focusing on TD learning with linear function approximation, polynomial-decaying stepsize, Polyak-Ruppert averaging and $i.i.d.$ samples, this paper makes the following theoretical and methodological contributions.
  1. 1.
    Improved high-probability guarantees. We show that, for an appropriate choice of the initial stepsize, the $L_{2}$ norm of the estimation error is of order
    $O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{d}{\delta}}\right),$
    with probability at least $1-\delta$, where $T$ denotes the number of iterations, $\bm{\Lambda}^{\star}$ represents the asymptotic variance matrix of the estimation error, and $\mathsf{Tr}(\bm{\Lambda}^{\star})$ its trace. To the best of our knowledge, this is the first high-probability convergence guarantee for this problem that matches the asymptotic variance, up to logarithm factors, without requiring a projection step to control the iterative trajectory. Table 1 compares our high-probability convergence guarantees with the latest results in the literature.
  2. 2.
    Sharper Berry-Esseen bounds. We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets, achieving a $O(T^{-\frac{1}{3}})$ rate as the function of the number of iterations. Our results improve upon the sharpest rate to date, or order $O(T^{-\frac{1}{4}})$, due to Samsonov et al., (2024).
  3. 3.
    Novel variance estimator. We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$, and establish a converge rate for $\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ of order $O(T^{-\frac{1}{3}})$ in Frobenius norm.
@@ 1.1 Contributions
We also verify our theoretical results by numerical simulations. As an intuitive illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (as shown in red curves).
1.2 Notation
For any complex number $z\in\mathbb{C}$, we use $\mathsf{Re}(z)$ to denote its real part. Throughout the paper, we use boldface small letters to denote vectors and capital letters to denote matrices. For any vector $\bm{x}$, we use $\|\bm{x}\|_{2}$ to denote its $L_{2}$ norm; for any matrix $\bm{M}$, we use $\mathsf{Tr}(\bm{M})$ to denote its trace, $\mathsf{det}(\bm{M})$ to denote its determinant, $\mathsf{cond}(\bm{M})$ to denote its conditional number, $\|\bm{M}\|$ to denote its spectral norm (i.e. the largest singular value), and $\|\bm{M}\|_{\mathsf{F}}$ to denote its Frobenius norm, i.e. $\|\bm{M}\|_{\mathsf{F}}=\sqrt{\mathsf{Tr}(\bm{M}^{\top}\bm{M})}$. For any symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(\bm{M})=\lambda_{1}(\bm{M})\geq\lambda_{2}(\bm{M})\geq...\geq\lambda_{d}(\bm{M})=\lambda_{\min}(\bm{M})$ to denote its eigenvalues. If a matrix $\bm{M}$ is positive (semi-)definite, i.e. $\lambda_{\min}(\bm{M})>(\geq)0$, we write $\bm{M}\succ(\succeq)\bm{0}$; when two matrices $\bm{X}$ and $\bm{Y}$ satisfy $\bm{X}-\bm{Y}\succ(\succeq)\bm{0}$, we write $\bm{X}\succ(\succeq)\bm{Y}$ or equivalently $\bm{Y}\prec(\preceq)\bm{X}$. For a vector $\bm{x}\in\mathbb{R}^{d}$, we use $\{x_{i}\}_{1\leq i\leq d}$ to denote its entries; for a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, we use $\{X_{ij}\}_{1\leq i\leq m,1\leq j\leq n}$ to denote its entries. The vectorization of matrix $\bm{X}\in\mathbb{R}^{m\times n}$, denoted as $\textbf{vec}(\bm{X})$, is defined as a vector in $\mathbb{R}^{mn}$ with entries $[\textbf{vec}(\bm{X})]_{(i-1)m+j}=\bm{X}_{ij}$ for every $1\leq i\leq m$ and $1\leq j\leq n$. For any matrices $\bm{X}\in\mathbb{R}^{m\times n}$, and $\bm{Y}\in\mathbb{R}^{p\times q}$, their Kronecker product, denoted as $\bm{X}\otimes\bm{Y}$, is a matrix of the size of $mp\times nq$, and can be written in blocked form as
With subscripts, it can be represented as $(\bm{X}\otimes\bm{Y})_{(i-1)m+j,(k-1)p+\ell}=X_{ik}Y_{j\ell}$.
For sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$ of numbers, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to rsignify that there exists a quantity $C>0$ independent of $t$ such that $f_{t}\leq Cg_{t}$ for all $t$. We will use the notatin $\widetilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters. We will measure the difference between two probability distributions on $\mathbb{R}^{d}$ using the convex distance, the maximal distance convex sets. Specifically, for any two probability distributions $P$ and $Q$, we set
$d_{\mathsf{C}}(P,Q)=\sup_{\mathcal{A}\in\mathscr{C}_{d}}|P(\mathcal{A})-Q(\mathcal{A})|.$
@@ 1.2 Notation
where $\mathscr{C}_{d}$ is the class of all convex set in $\mathbb{R}^{d}$. We use $d_{\mathsf{TV}(P,Q)}$ to denote the Total Variation (TV) distance between $P$ and $Q$.
paper stepsize convergence guarantee constraint on $t,T,\delta$
[…]
Li et al., 2024b $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\widetilde{C}$
Samsonov et al., (2023) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\widetilde{C}$
This work, Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{d}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\widetilde{C}$
Table 1: Comparisons between our results on convergence rates for TD with prior work that derived with high probability control. The quantity $\widetilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
@@ 2 Background
An infinite-horizon Markov Decision Process (MDP) (Bertsekas,, 2017; Sutton and Barto,, 2018) can be represented mathematically by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$. Here, $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent respectively, both assumed finite throughout this paper; the function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$; the reward function $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ measures the immediate reward to the agent after taking action $a$ at state $s$; finally, $\gamma\in(0,1]$ is the discount factor, used to assess the impact of future rewards.
In choosing their actions, the agent follows a randomized action selection rule, or policy, represented as a (time-invariant) conditional probability distribution $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ over the action space given the state of the environment. For any policy $\pi$, its associated reward function $r^{\pi}:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e.
$r^{\pi}(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],\quad s\in\mathcal{S}.$
For readability, we will omit the superscript $\pi$ and use $r(\cdot)$ to denote $r^{\pi}(\cdot)$. Any policy $\pi$ induces a transition kernel $P^{\pi}:\mathcal{S}\mapsto\Delta(\mathcal{S})$, representing the probability distribution of the next state, given the current state $s$ and the policy $\pi$. Formally,
$P^{\pi}(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a),\quad s^{\prime}\in\mathcal{S}.$
We assume throughout that there exists a unique stationary distribution $\mu^{\pi}\in\Delta(\mathcal{S})$ for the transition matrix ${P}^{\pi}$, assigning positive probability to each state. That is,
$0<\mu^{\pi}(i)=\mathbb{P}_{s\sim\mu^{\pi},s^{\prime}\sim P^{\pi}(\cdot\mid s)}(s^{\prime}=i),\quad\forall i\in\mathcal{S}.$
@@
The worth of a policy $\pi$ is measured through its value function $V^{\pi}:\mathcal{S}\mapsto\mathbb{R}_{+}$ returning, for each initial state, the expected value of all cumulative future rewards, appropriately discounted, when policy $\pi$ is executed under the MDP $\mathcal{M}$. Formally,
$V^{\pi}(s):=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\,\big{|}\,s_{0}=s\right],\quad s\in\mathcal{S}.$ (1)
A fundamental, defining feature of the value function is that it satisfies a certain self-consistency property, known as the Bellman equation, a functional fixed-point equation of the form
$V^{\pi}(s)=r(s)+\mathbb{E}_{s^{\prime}\sim P^{\pi}(\cdot\mid s)}V^{\pi}(s^{\prime})=:\mathcal{T}V^{\pi}(s),$ (2)
where $\mathcal{T}$ is the Bellman’s operator. Indeed, the value function is the unique solution to the Bellman equation. With a slight abuse of notation, we shall omit the superscript $\pi$ in $V^{\pi}$ and abbreviate it as $V$ as the target policy is considered fixed.
@@
In most practical scenarios, though finite, the state space $\mathcal{S}$ is prohibitively large. As a result, the value function is often very complex and impractical to compute. To overcome this issue, it is customary to consider a parametric approximation of $V$ through a linear function, which is computationally tractable and amenable to statistical estimation (see, e.g. Tsitsiklis and Van Roy, (1997); Dalal et al., (2018); Bhandari et al., (2021); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018)) In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a feature vector in $\mathbb{R}^{d}$, where $d$ is usually smaller than $|\mathcal{S}|$, the value function at $s$, $V(s)$, is expproximated by
$V_{\bm{\theta}}(s):={\bm{\phi}}(s)^{\top}{\bm{\theta}},\qquad\quad s\in\mathcal{S},$ (3)
where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients.
Computing the optimal linear approximation to $V$ can be cast as a least-squares minimization problem over $\mathbb{R}^{d}$ of finding a coefficient vector $\bm{\theta}$ minimizing the expected difference between $V_{\bm{\theta}}$ and its Bellman projection $\mathcal{T}V_{\bm{\theta}}$. That is,
$\bm{\theta}^{\star}:=\underset{\bm{\theta}\in\mathbb{R}^{d}}{\text{argmin }}\mathbb{E}_{s\sim\mu}(V_{\bm{\theta}}(s)-\mathcal{T}V_{\bm{\theta}}(s))^{2},$ (4)
where we recall that $\mu$ that denote the stationary distribution of $P^{\pi}$. It turns out that the problem (4) admits a unique solution (Tsitsiklis and Van Roy,, 1997), given by
@@
$\bm{A}$ $:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d},$ (6a)

### v3

# Statistical Inference for Policy Evaluation with Temporal Difference Learning

@@ Abstract
We investigate the statistical properties of Temporal Difference (TD) learning with Polyak-Ruppert averaging, arguably one of the most widely used algorithms in reinforcement learning, for the task of estimating the parameters of the optimal linear approximation to the value function. We make three significant contributions that improve the current state-of-the-art results: (i) we derive sharper high probability convergence guarantee that depend explicitly on the asymptotic variance and hold under weaker conditions than those normally assumed; (ii) we establish refined high-dimensional Berry-Esseen bounds over the class of convex sets, achieving faster rates than those previously established in the literature, and (iii) we propose and analyze a novel, computationally efficient online plug-in estimator of the asymptotic covariance matrix. These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function approximation, with guaranteed finite-sample coverage. We demonstrate the applicability of our theoretical findings through numerical experiments.
@@ 1 Introduction
The feasibility and accuracy of statistical inference tasks, such as hypothesis testing or constructing confidence regions or simultaneous confidence intervals, are traditionally predicated on using estimators that are not only consistent but also satisfy some type of distributional approximation. In most cases, this is accomplished via Central Limit Theorems (CLTs), which ensure that the procedures in use admit a Gaussian approximation. Due to the high dimensionality of modern scientific and engineering applications, there has been a surge of interest in establishing new high-dimensional Berry-Esseen bounds, CLTs that hold in high dimensions and have a finite-sample guarantee. While efforts in this area have focused on Berry-Esseen bounds for normalized sums, few results are available to handle more complicated scenarios, where estimators are computed in a nonlinear, iterative manner. In this paper, we investigate one such scenario and study in detail the consistency, Gaussian approximations, and statistical inference for the output of the Temporal Difference (TD) learning algorithm for policy evaluation with linear function approximation in Reinforcement Learning (RL).
A critical task in RL is to evaluate with confidence the quality of a given policy, measured by its value function. As RL models are based on Markov Decision Processes (MDPs), statistical inference must accommodate the online nature of the sampling mechanism. TD learning (see, e.g., Sutton, 1988), an iterative procedure that can be framed as an instance of stochastic approximation (SA), is arguably the most widely used and effective algorithm for policy evaluation in RL. The last several years have witnessed a flurry of activities within the statistical machine learning community to derive finite sample consistency rates of TD iterates, leading to deep results and a much improved understanding of the performance and limitations TD learning. See, e.g., Khamaru et al. (2020); Li et al. (2024); Bhandari et al. (2021); Dalal et al. (2018); Srikant and Ying (2019); Lakshminarayanan and Szepesvari (2018); Mou et al. (2020); Durmus et al. (2022); Li et al. (2024); Samsonov et al. (2024); Prashanth et al. (2013); Borkar et al. (2025) for a partial and growing list of relevant references. Though for the tabular case (finite state and action spaces) the TD algorithm is known to be minimax rate optimal Li et al. (2024), its properties – and, in particular, its sample complexity – in more complex and high-dimensional settings have not been fully understood.
paper stepsize convergence guarantee constraints on $t,T,\delta$
[…]
Li et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\tilde{C}$
Samsonov et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\tilde{C}$
Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{d}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\tilde{C}$
Table 1: Comparisons between our results on TD convergence rate with prior work. The quantity $\tilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
In a more specialized line of work, concerned with establishing distributional approximations of the TD iterates, it has been shown that, using Polyak-Ruppert averaging (Polyak and Juditsky, 1992; Ruppert, 1998), the TD estimation error converges to a Gaussian distribution in the large sample limit; see, e.g., Fort (2015); Mou et al. (2020); Li et al. (2023). While these results provide a formal justification for the validity of statistical inference, the guarantees provided in those contributions are non-quantitative, in the sense that they are of asymptotic nature and hold only in fixed dimensions. As a result, they do not adequately capture the dependence on the sample size and other problem-related quantities, most notably the dimensionality, all of which affect the accuracy of statistical inference. In this paper, we partially address this gap. We focus on TD learning with Polyak-Ruppert averaging for estimating the coefficients of the best linear approximation of the value function, assuming i.i.d. samples. We make the following theoretical and methodological contributions.
  1. 1.
    Improved high-probability guarantees. We show that, for an appropriate choice of the initial stepsize, the Euclidean norm of the estimation error for the TD learning algorithm is of order
    $O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}\right),$
    with probability at least $1-\delta$, where $T$ denotes the number of iterations, $\bm{\Lambda}^{\star}$ represents the asymptotic variance matrix of the estimation error, and $\mathsf{Tr}(\bm{\Lambda}^{\star})$ its trace. To the best of our knowledge, this is the first high-probability convergence guarantee for this problem that matches the asymptotic variance, up to logarithm factors, without requiring a projection step to control the iterative trajectory. Furthermore, our assumptions are weaker than those used in the literature. Table 1 compares our high-probability convergence guarantees with the latest available results.
  2. 2.
    Sharper Berry-Esseen bounds. We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets and using an estimated variance, achieving an $O(T^{-\frac{1}{3}})$ rate as a function of the number of iterations, which appears to be the fastest in the literature. While completing this work, we became aware of the recent related paper by Samsonov et al. (2024), which, under the same settings, derived a high-dimensional Berry-Esseen bound for linear SA with Polyak-Ruppert averaging, obtaining a slower rate, of order $O(T^{-\frac{1}{4}})$. See Section 3.2 for detailed comparisons.
  3. 3.
    Novel, computationally efficient variance estimator. We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$. We also show that the total variation distance between the (computable) distribution $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ vanishes with high probability at a rate $O(T^{-\frac{1}{3}})$.
@@
We verify our theoretical findings with some numerical simulations, discussed in the appendix. As an illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (shown by the red curves).
The i.i.d. condition. Our results are derived under the simplifying but standard assumption of i.i.d. data, which has been widely used in the recent literature on statistical analyses of TD learning, SA and stochastic gradient descent (see, e.g. Sheshukova et al., 2025). Analogous results, obtained using different techniques and yielding slower rates, have been recently obtained for the more challenging case of Markovian data: see Srikant (2024) and Wu et al. (2025).
Notation. We use boldface small letters to denote vectors and capital letters to denote matrices. For a vector $\bm{x}$, $\|\bm{x}\|_{2}$ incates its Euclidean norm; for a matrix $\bm{M}$, we use $\mathsf{Tr}(\bm{M})$, $\mathsf{det}(\bm{M})$, $\|\bm{M}\|_{\mathsf{F}}$ and $\|\bm{M}\|$ to denote its trace, determinant, Frobenius and spectral norm, respectively. For a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(\bm{M})=\lambda_{1}(\bm{M})\geq\ldots\lambda_{d}(\bm{M})=\lambda_{\min}(\bm{M})$ to denote its eigenvalues and $\mathrm{cond}(\bm{M})$ to indicate its condition number. For a vector $\bm{x}\in\mathbb{R}^{d}$ and a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we let $\|\bm{x}\|_{\bm{M}}=\sqrt{\bm{x}^{\top}\bm{Mx}}$. We will use the notation $\tilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters. Similarly, for positive sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to signify that $f_{t}\leq\tilde{C}g_{t}$, for some $C>0$ independent of $t$. The convex distance between two probability distributions on $\mathbb{R}^{d}$, say $P$ and $Q$, is defined as
$d_{\mathsf{C}}(P,Q)=\sup_{\mathcal{A}\in\mathscr{C}_{d}}|P(\mathcal{A})-Q(\mathcal{A})|,$
@@
where $\mathscr{C}_{d}$ is the class of all convex subsets of $\mathbb{R}^{d}$. We use $d_{\mathsf{TV}}(P,Q)$ to denote instead the total variation (TV) distance between $P$ and $Q$. Notice that $d_{\mathsf{C}}(P,Q)\leq d_{\mathsf{TV}}(P,Q)$.
@@ 2 Background
An infinite-horizon Markov Decision Process (MDP) (Bertsekas, 2017; Sutton and Barto, 2018) is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent, respectively, both assumed for simplicity to be finite (though our results hold for countable spaces). The function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$. The reward function $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ measures the immediate reward to the agent after taking action $a$ at state $s$; finally, $\gamma\in[0,1)$ is the discount factor, used to assess the impact of future rewards.
In choosing their actions, the agent follows a randomized action selection rule, or policy, represented as a (time-invariant) conditional probability distribution $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ over the action space given the state of the environment. For any policy $\pi$, its associated reward function $r:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e. $r(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],$ for $s\in\mathcal{S}.$ Any policy $\pi$ induces a transition kernel $P$, representing the probability distribution of the next state given the current state $s$ and the policy $\pi$, given by $P(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a)$ for $s,s^{\prime}\in\mathcal{S}.$ We assume throughout that there exists a unique stationary distribution $\mu$ for the transition matrix ${P}$; that is, $\mu(i)=\mathbb{P}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}(s^{\prime}=i)$, $i\in\mathcal{S}$.
@@
The worth of a policy $\pi$ is measured by its value function $s\in\mathcal{S}\mapsto V(s):=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\,\big{|}\,s_{0}=s\right]$ returning, for each initial state, the expected value of all cumulative future rewards, appropriately discounted. A fundamental, defining feature of the value function is that it arises as the unique solution to the the Bellman equation, a functional fixed-point equation of the form
$V(s)=r(s)+\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}V(s^{\prime})=:\mathcal{T}V(s).$ (1)
where $\mathcal{T}$ is known as Bellman’s operator.
@@
In most practical scenarios, the value function is very complex and impossible to compute due to the prohibitively large size of the state space $\mathcal{S}$ A common way to overcome this issue is to consider a parametric approximation of $V$ through a linear function, which is both computationally tractable and amenable to statistical estimation. In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a centered feature vector in $\mathbb{R}^{d}$, where $d$ is typically small(er than $|\mathcal{S}|$), a linear value function approximation is given by
$s\in\mathcal{S}\mapsto V_{\bm{\theta}}(s):={\bm{\phi}}(s)^{\top}{\bm{\theta}},$ (2)
where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients. Throughout this paper, we assume that the Eculidean norm of the features are uniformly bounded, i.e., $\|\bm{\phi}(s)\|_{2}\leq 1$ for all $s\in\mathcal{S}$.
The best approximating linear function is the one parametrized by the coefficient $\bm{\theta}^{\star}$ solving the projected Bellman equation (Tsitsiklis and Van Roy, 1997) $V_{\bm{\theta}^{\star}}=\Pi\mathcal{T}V_{\bm{\theta}^{\star}},$ with $\Pi$ the orthogonal projection onto the subspace of functions that can be represented by linear combinations of the features. See, e.g., Duan et al. (2021). In turn, this (unique!) solution is given by
@@
$\bm{A}:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d}\quad\text{and}\quad\bm{b}:=\mathop{\mathbb{E}}\limits_{s\sim\mu}\left[\bm{\phi}(s)r(s)\right]\in\mathbb{R}^{d}.$ (4)
Throughout, we let
$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)],\quad\lambda_{0}:=\lambda_{\min}(\bm{\Sigma}),\quad\text{and}\quad\lambda_{\Sigma}:=\lambda_{\max}(\bm{\Sigma})$ (5)
be the feature covariance matrix and its smallest and largest eigenvalues, respectively.
The Temporal Difference (TD) learning with Polyak-Ruppert averaging.
While in most real-world RL scenarios the state-action-reward tuples follow a Markov decision process, many of the most recent theoretical contributions – including those delivering the sharpest guarantees listed in Table 1 and Samsonov et al. (2024) – have made the simplifying but technically very convenient assumption of $i.i.d.$ data. Here we follow the same convention and analyze the properties of the TD learning algorithm applied to a sequence of $T$ $i.i.d.$ state-action-reward tuples of the form
$\{(s_{t},a_{t},s_{t}^{\prime},r_{t})\}_{t=1}^{T},$ (6)
where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$, the stationary distribution, and, for eact $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$.
The TD learning algorithm is a stochastic approximation method to find in an iterative manner the solution (3). In detail, for a sequence of pre-selected step sizes $\{\eta_{t}\}_{t\geq 1},$ TD proceeds using the update rule
$\bm{\theta}_{t}=\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}),$ (7)
where
$\bm{A}_{t}:=\bm{\phi}(s_{t})\left(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})\right)^{\top}\quad\text{and }\quad\bm{b}_{t}:=r_{t}\bm{\phi}(s_{t}).$ (8)
Rather than the final TD iterate $\bm{\theta}_{T}$, to estimate $\bm{\theta}^{\star}$ it is often preferred to use instead the average of all the iterates
$\overline{\bm{\theta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\bm{\theta}_{t}.$ (9)
This estimator is an instance of the Polyak-Ruppert averaging scheme (Polyak and Juditsky, 1992; Ruppert, 1998), a stochastic approximation methodology that provably reduces the variability and accelerates convergence. Furthermore, with a proper choice of the stepsizes, the averaged iterates converge to a Gaussian distribution, thus enabling statistical inference. To date, the best Berry-Esseen bounds rates for stochastic approximation methods and TD learning with Polyak-Ruppert averaging are due to Samsonov et al. (2024), which in turn are based on general non-linear Berry-Esseeen bounds by Shao and Zhang (2022). In this paper we improve on these results by deriving novel and sharper high-probability and Berry-Esseen bounds for the estimation error of the TD algorithm
We conclude by mentioning that the choice of learning rates $\{\eta_{t}\}_{t\geq 0}$ in the updates (7) plays a critical role in determining the statistical properties of (both the original and) the averaged TD estimator (9). Here we assume a polynomial-decaying stepsizes schedule $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(1/2,1)$ and $\eta_{0}<1/2$.
3 Results
It is known (Polyak and Juditsky, 1992; Fort, 2015) that the TD estimation error of (10) satisfies